40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jul 26 22:45:14 2013, manasa, Update, General, Vent progress 
    Reply  Tue Jul 30 00:21:15 2013, Koji, Update, General, Vent progress 
       Reply  Tue Jul 30 11:30:55 2013, manasa, Update, General, ITMX oplev 
          Reply  Tue Jul 30 13:41:35 2013, Koji, Update, General, ITMX oplev 
             Reply  Tue Jul 30 15:38:24 2013, manasa, Update, General, ITMX oplev 
Message ID: 8937     Entry time: Tue Jul 30 11:30:55 2013     In reply to: 8936     Reply to this: 8938
Author: manasa 
Type: Update 
Category: General 
Subject: ITMX oplev 

Quote:

[Koji, Manasa, Sujan]

The ITMX oplev was also checked from the windows without removing the door.
The beam is actually hitting the right positions of the steering mirror and the test mass
although there are lots of stray beams because of the BS oplev and some halo assciated with the ITMX He-Ne laser(!?).

[Steve, Gautam, Manasa]

While we checked the ITMX oplev situation yesterday, we found that the beam hitting the ITM and the in-vacuum steering mirrors had  a halo around them. We used the set of irides in the path of the ingoing beam and cut the stray light around the beam. This reduced the intensity of the halo around the mirrors. We noticed that the halo accounted for 2000 counts of 6500 at the oplev QPD. We tried changing the laser and this did not make the situation any better.

Also there are a couple or more strong stray beams from the BS oplev.

Thoughts:
I suspect that the BS oplev leakage is messing with the ITMX oplev. Why?? We have been seeing the breathing of transmitted beam from the X arm cavity and the shadow sensor readings have been bigger than usual since the last vent (when we changed the BS oplev path). Also, the ingoing oplev beam is close to clipping at the PR2 stack.

I think it would be best to open the ITMX chamber and modify the in-vac steering mirror layout of the oplev.

ELOG V3.1.3-